### v1

Abstract
Context. Gamma-ray bursts (GRBs), observed at redshifts as high as 9.4, could serve as valuable probes for investigating the distant Universe. However, this necessitates an increase in the number of GRBs with determined redshifts, as currently, only 12% of GRBs have known redshifts due to observational biases.
Aims. We aim to address the shortage of GRBs with measured redshifts, enabling us to fully realize their potential as valuable cosmological probes
Methods. Following Dainotti et al. (2024c), we have taken a second step to overcome this issue by adding 30 more GRBs to our ensemble supervised machine learning training sample, an increase of 20%, which will help us obtain better redshift estimates. In addition, we have built a freely accessible and user-friendly web app that infers the redshift of long GRBs (LGRBs) with plateau emission using our machine learning model. The web app is the first of its kind for such a study and will allow the community to obtain redshift estimates by entering the GRB parameters in the app.
Results. Through our machine learning model, we have successfully estimated redshifts for 276 LGRBs using X-ray afterglow parameters detected by the Neil Gehrels Swift Observatory and increased the sample of LGRBs with known redshifts by 110%. We also perform Monte Carlo simulations to demonstrate the future applicability of this research.
Conclusions. The results presented in this research will enable the community to increase the sample of GRBs with known redshift estimates. This can help address many outstanding issues, such as GRB formation rate, luminosity function, and the true nature of low-luminosity GRBs, and enable the application of GRBs as standard candles.
Key Words.:
gamma-ray bursts – machine learning – redshift
1 Introduction
Gamma-Ray Bursts (GRBs) are the brightest and most powerful explosion events after the Big Bang and are observed across a wide range of redshifts, from 0.0085 (Galama et al., 1998) to 8.2 and 9.4 (Tanvir et al., 2008; Cucchiara et al., 2011). This vast range of redshifts ($z$) makes GRBs ideal probes for studying the early Universe and tracking its evolution, providing insights into the composition, expansion rate, and other key aspects. GRBs, generally detected in $\gamma$-rays, X-rays, and occasionally in optical wavelengths, are classified into two types based on their duration, characterized by the $T_{90}$ parameter. $T_{90}$ is the time span during which the GRB emits 90% of its total observed fluence in $\gamma$-rays. GRBs with $T_{90}<2s$ or $T_{90}>2s$ and are classified as short GRBs (SGRBs) and long GRBs (LGRBs), respectively.
LGRBs are produced from core collapse supernovae (Mazets et al., 1981; Kouveliotou et al., 1993; Paczyński, 1998; Woosley & Bloom, 2006), while SGRBs are produced by neutron star-neutron star or neutron star-black hole mergers (Narayan et al., 1992; Duncan & Thompson, 1992; Usov, 1992; Thompson, 1994; Nakar, 2007; Goldstein et al., 2017a, b; Abbott et al., 2017).
In addition to these two main categories of GRBs, numerous other sub-classes have been identified in the literature. These classes are based only on observational properties such as the $T_{90}$ duration, spectral, and lightcurve properties. These include classes such as Short GRBs with extended emission (SEE, Norris et al. (2000); Norris & Bonnell (2006); Levan et al. (2007); Norris et al. (2010)), where $T_{90}>2$ seconds but with similar spectral properties to SGRBs, Very Long GRBs (VLGRBs, Levan et al. (2014)) and Ultra Long GRBs (ULGRBs, Stratta et al. (2013); Zhang et al. (2014); Gendre et al. (2019) where $T_{90}>500$ and $1000$ seconds, respectively. Zhang et al. (2007) unified these various classes based on their GRB progenitors. According to their prescription, Type-I GRBs comprise of classes which are SGRBs, SEE, and others similar to SGRBs, and occur due to compact binary mergers (Nakar, 2007; Abbott et al., 2017). Type-II GRBs comprise LGRBs, VLGRBs, ULGRBs, and other similar classes with core-collapse supernova progenitors (the so-called collapsar events) (Woosley, 1993; Woosley et al., 1993). Irrespective of these efforts, an exhaustive classification of GRBs has not yet been found, and this is an active field of research.
Phenomenologically, the GRB lightcurve (LC) from across all the classes can be separated into two distinct phases. The prompt emission is the main emission, characterized by a burst of high-energy photons spanning from $\gamma$-rays to hard X-rays and sometimes extending into optical wavelengths (Vestrand et al., 2005; Beskin et al., 2010; Gorbovskoy et al., 2012; Vestrand et al., 2014). Following the prompt, the afterglow is a long-lasting, multiwavelength emission powered by the interaction of the GRB ejecta with the surrounding medium (van Paradijs et al., 1997; Costa et al., 1997; Piro et al., 1998). It is detected across various wavelengths, such as X-rays, optical, and occasionally even radio.
A peculiar trait of GRBs, uncovered by the Neil Gehrels Swift observatory (hereafter Swift, Gehrels et al. (2004)), is the presence of a flat region in the LC, the so-called plateau emission, which is characterized by almost constant flux values (Nousek et al., 2006; Zhang et al., 2006; O’Brien et al., 2006; Sakamoto et al., 2007; Liang et al., 2007b; Dainotti et al., 2008; Zaninoni et al., 2013; Rowlinson et al., 2014a). The end of the plateau is considered to be a point in the LC’s afterglow, where the luminosity begins to decay with time as a powerlaw. The plateau feature has been found in 42% of X-ray afterglows (Evans et al., 2009; Li et al., 2018). In our recent estimate, the percentage of GRBs with observable plateaus is 38.5%.
Currently, the primary hurdle in the LGRB population study is the lack of a large GRB sample with known redshift. Quick localization and spectral data are required to determine GRBs’ redshift.
Swift allows for this quick detection and multi-wavelength follow-up in $\gamma$-rays, X-rays, and ultraviolet/optical. It has three telescopes: the Burst Alert Telescope (BAT, Burrows et al. (2005)), the X-ray Telescope (XRT, Barthelmy et al. (2005)), and the Ultraviolet/Optical Telescope (UVOT, Roming et al. (2005)). Swift uses XRT for precise localization, and with UVOT, it obtains spectra and sometimes redshifts. Swift’s rapid localization abilities have significantly contributed to exploring the high-$z$ Universe.
However, although Swift triggers many observation programs, only 26% (431) of Swift’s 1656 GRBs have known spectroscopic redshifts (as of 20th August 2024). The Fermi Telescope has observed more than 3700 GRBs, with its Large Area Telescope (LAT, Ajello et al. (2019)) and the Fermi-Gamma-Ray Burst Monitor (GBM, Meegan et al. (2009)). Currently, only 12% of GRBs detected by Swift and Fermi have spectroscopic redshift ($z_{obs}$), $\sim$90of which are observed by Swift.
Measuring redshift, especially high-$z$, is challenging due to limited telescope time and the difficulty of securing time for follow-up observations using large aperture telescopes. Therefore, determining the $z$ of GRBs via alternative methods is crucial.
Correctly determining the GRB luminosity function (LF) is one of the main goals that can be achieved by increasing the number of GRBs with known redshifts. The LF gives us the number of GRBs per unit luminosity and is vital for understanding the properties of the GRB population, such as their energy release and emission mechanisms. Another principal objective is determining the cosmic GRB formation rate (GRBFR). The GRBFR measures the frequency of GRBs across cosmic time.
Petrosian et al. (2015); Petrosian & Dainotti (2024) have found a discrepancy at low redshift ($10<z<$ 1) between the LGRB formation rate and the star formation rate (SFR). This discrepancy, which has been observed by several research groups independently, remains an active area of investigation. Addressing this requires a bigger sample of GRBs with known redshift, thus highlighting the importance of our study.
The correct determination of the GRB rate for both LGRBs and SGRBs is hampered by the broad range of GRB physical features; therefore, a precise distribution of these features and their evolution over cosmic time is necessary. Some of us investigated this issue with a sample of 179 GRBs with optical LCs (Dainotti et al., 2024b), where we determined the luminosity and formation rate evolutions and the general shape of the LF after correcting for evolution. We observed the average LGRB’s rate at low-$z$ ($z<1$) to be 10.46 Gpc-3 $yr^{-1}$, which is roughly 3 times more than Hopkins & Beacom (2006) SFR (3.7 Gpc-3 $yr^{-1}$). This LGRB rate is an actual rate where the normalization is in physical units, which contrasts with the arbitrary normalization used in most current literature.
To address the challenges outlined above, we propose here a machine learning (ML) based approach for estimating the redshift of GRBs. This work is a follow-up of Dainotti et al. (2024c) and Dainotti et al. (2024b) and utilizes an enlarged sample, 20% larger than the sample in Dainotti et al. (2024c), to train our ML models. In addition, we have extended the generalization sample by 35%. This is the set of GRBs without known redshifts, and we predict the redshift of 276 GRBs using our trained ML model. Furthermore, we release a user-friendly web application that provides redshift estimates for new GRBs, built on the ML models described in this paper.
The paper is structured as follows: Sec. 2 discusses the advantages of using an ML approach compared to using linear relations for the redshift estimation. In Sec. 3, we describe our dataset. In Sec. 4, we describe the functionalities and various modules of our web app, as well as details of the ML models being used. We also describe the Monte Carlo (MC) simulation performed to understand the viability of our ML models in relation to future data. Sec. 5 shows the results obtained from our analysis and the usability of the web app. Here, we also present the future applicability of our ML approach. In Sec. 6, we discuss the relevance of our results. Finally, we summarize and conclude in Sec. 7.
2 Advantage of machine learning VS linear relationships
Expanding the GRB sample with redshift provides a significant benefit, notably the possibility of using them as standardized candles, objects with known intrinsic luminosities, or with luminosities that can be derived via established relationships among GRB properties. The adoption of empirical relations between the distance-dependent and intrinsic properties of GRBs enables cosmological studies in unexplored redshift ranges. One of the earliest efforts is the Dainotti Relation (Dainotti et al., 2008, 2011, 2015, 2017), a roughly inversely proportional relationship between the rest-frame time at the end of the plateau phase ($T_{a}/(1+z)$) and its corresponding X-ray luminosity ($L_{a}$). Later, Dainotti et al. (2013) showed via the use of the Efron and Petrosian method (Efron & Petrosian (1992)), this relation is intrinsic and not due to selection biases nor the redshift evolution. The Dainotti Relation has also been observed in radio wavelengths (Dainotti et al., 2020; Levine et al., 2022). It has also been extended in three dimensions in X-ray, optical (Dainotti et al., 2022a), and $\gamma$-rays (Dainotti et al., 2021b), where the peak prompt luminosity ($L_{peak}$) has been added to the two-dimensional Dainotti relation (Dainotti et al., 2016, 2017, 2020).
The Dainotti relationships (both in two and three dimensions) were a subject of cosmological investigations, and have been utilized as valuable tools (Cardone et al., 2009, 2010; Dainotti et al., 2013; Postnikov et al., 2014; Cao et al., 2021, 2022b, 2022a; Dainotti et al., 2022b, 2023; Bargiacchi et al., 2023; Dainotti et al., 2023). The Dainotti relations have been shown to provide compatible results for $\Omega_{M}$ (the matter density parameter) under the assumption of flat $\Lambda$CDM model when used in conjunction with supernovae (SNe) Ia (Dainotti et al., 2022b). This approach also has the advantage of expanding the distance ladder up to $z$ = 5, which is much higher than the most distant SNe Ia observed up to date ($z$ = 2.9) (Pierel et al., 2024). Wang et al. (2022) further extended the Dainotti relation up to $z=5.91$. They also standardized the LGRB sample using a method similar to the Phillips Correlation (Phillips, 1993). This standardization helped Wang et al. (2022) achieve a lower scatter on the rescaled luminosity, thus highlighting the importance of the plateau phase as a possible standard candle.
Another notable effort in establishing a fundamental GRB relation is the so-called Combo relation (Izzo et al., 2015), joining $L_{a}$, $T_{a}/(1+z)$, powerlaw decay of afterglow, and intrinsic spectral peak energy of prompt (Ep,i). Izzo et al. (2015) demonstrated, with a limited sample of 5 GRBs, a minimization of the dependency on SNe Ia for calibrating GRBs. They use a two-step calibration process for the Combo relation and measured $\Omega_{M}=0.29^{+0.23}_{-0.15}$ in $\Lambda$CDM.
Recently, Yi et al. (2022) analyzed samples of 174 X-ray LCs of GRBs with plateau and 104 GRBs with flares, overlapping by 51 sources. They discovered three tight correlations between plateau energy ($E_{plat}$), flare energy ($E_{flare}$) and isotropic prompt energy ($E_{\gamma,iso}$). The correlations were between $E_{plat}-E_{\gamma,iso}$, $E_{flare}-E_{\gamma,iso}$ and $(E_{plat}+E_{flare})-E_{\gamma,iso}$. $E_{plat}-E_{\gamma,iso}$ correlation is similar to the $L_{a}-L_{peak}$ Dainotti correlation (Dainotti et al., 2011), thus independently confirming it. Yi et al. (2022) also conclude that X-ray plateaus and flares may originate in the same way but, due to the effects of the surrounding medium, appear as different features in the afterglow LC.
The extension of the Dainotti relations in three dimensions has also been achieved by Xu & Huang (2012), where they add the isotropic $\gamma$-ray energy to the two-dimensional relation, naming it the $L-T-E$ correlation. Zhao et al. (2019); Tang et al. (2019) independently confirmed the existence of this correlation for 174 GRBs from Swift. Tang et al. (2019) also improved the best-fit relation to be $L_{a}\propto T_{a}^{-1.01}E^{0.84}$. This work was expanded upon further by Xu et al. (2021) and Deng et al. (2023). Xu et al. (2021) showed the redshift evolution of the $L-T-E$ correlation and introduced a new correlation with the spectral peak energy labeled $L-T-E_{p}$ correlation. Deng et al. (2023) updated coefficients of the $L-T-E$ correlation of Tang et al. (2019) to $L_{a}\propto T_{a}^{-0.99}E^{0.86}$ with 210 LGRBs with plateau and also derived pseudo-redshifts for 108 plateau GRBs. However, their redshift estimates are limited by intrinsic scatter and functional form of the $L-T-E$ correlation. Moreover, the precision of their approach is not tested against GRBs with known redshift. In comparison, our ML-based approach tests the performance of our model on events with known redshifts.
Previously, there have been several other efforts aiming to enhance the count of GRBs with known redshifts by exploring correlations between distance-independent parameters (such as peak flux and afterglow plateau duration) and distance-dependent GRB characteristics (such as prompt emission and peak luminosity) to estimate pseudo-redshifts for GRBs lacking observed values(Reichart, 2001; Atteia, 2003, 2005; Yonetoku et al., 2004; Dainotti et al., 2011). Only very few GRBs can reach the small uncertainty (5%) (Guiriec et al., 2016). Specifically, when using a relation between the rest-frame $T_{a}$ and its luminosity, only 28% of cases have uncertainties $\Delta z/z_{obs}<0.90$ (Dainotti et al., 2011). If the redshift is extracted by employing the Amati correlation (Amati et al., 2002) between $E_{peak}$ (the peak in the $\nu F_{\nu}$ spectrum) and the energy emitted isotropically during the prompt emission, the Pearson coefficient ($r$) of correlation between $z_{obs}$ and predicted redshift ($z_{pred}$) is $0.67$. Amati (2006) provided $z_{pred}$ for 17 GRBs, and these are accurate within a factor of 2.
Another issue, which is often overlooked, is that these pseudo-redshifts are inherently subject to inaccuracies. This is because of their non-linear relation to the luminosity distance ($D_{L}$).
The major problem in the forward-fitting relationships is the dependence on cosmology. Indeed, one of the parameters of this relationship depends on the $D_{L}$, which itself depends on cosmological parameters. Thus, using any of these methods would induce a circularity problem, which must be avoided, especially when we tackle cosmology. In addition, given the functional form of $\log(D_{L})$, as it flattens at $z\gtrsim 3$, any small variation of $\log(D_{L})$ can become a large variation in redshift determination. Hence, the difficulty in determining the redshifts accurately with these forward fitting methods. Furthermore, based on relationships, the forward fitting methods mentioned do not account for the errors on the parameters, thus underestimating the realistic uncertainties on the redshifts. Thus, we have followed a route via ML algorithms to obtain more accurate predictions and overcome these issues.
Similar work has been done by Ukwatta et al. (2016); Aldowma & Razzaque (2024). However, here, we use the plateau features, whose correlations have less intrinsic scatter than prompt ones. This can ensure more stable ML results.
3 The Data Sample
Due to the diverse nature of various GRB classes, it is essential to avoid mixing the characteristics of these classes in an ML analysis. In a previous paper, some of us (Dainotti et al., 2024c) have pointed out that the removal of some additional classes, especially the SEE, does not significantly affect the prediction. Thus, this study focuses only on LGRBs, the most numerous and homogeneous subset. Thus, all SGRBs are removed from our sample. Our data comprises of LGRBs observed in $\gamma$-rays and X-rays by the BAT and XRT telescopes onboard the Swift. They are found in the Third Swift-BAT GRB Catalogue (Lien et al., 2016), and for data from 2016 until 12th December 2023, we use the NASA Swift GRB Search Tool. This sample is an extension of the one presented in Dainotti et al. (2024c), which itself was taken from Dainotti et al. (2020) and Srinivasaragavan et al. (2020). For training our models, we have removed the GRBs that have photometric redshifts; an example is the famous high-$z$ GRB090429B. Each GRB has ten properties (features). Four features belong to the prompt: the energy fluence over $T_{90}$ of the prompt emission in units of erg cm-2 (Fluence), $T_{90}$, the prompt peak photon flux in units of the number of photons cm-2 s-1 (Peak Flux), and from the BAT Telescope, the prompt photon index of the photon energy distribution modeled with a powerlaw (PhotonIndex). Six features belong to the X-ray plateau and afterglow: the time at the end of the plateau emission ($T_{a}$), its corresponding flux ($F_{a}$), the temporal powerlaw index after the end of the plateau emission ($\alpha$), the spectral index of the plateau ($\beta$) assuming a powerlaw for the spectral energy distribution, the neutral hydrogen column density along the line of sight (NH), the spectral index obtained as the time-averaged spectral fit from the Swift XRT Photon counting mode data ($\gamma$). The plateau emission feature is modeled using the so-called Willingale function (W07, Willingale et al. (2007)).
Some of the parameters such as the $Peakflux$, $T_{a}$, $F_{a}$, and $T_{90}$, are shown the Fig. 1, with $T_{a}$, $F_{a}$ representing the x and y coordinates of the black dot. All parameters of the afterglow were obtained in 0.3-10 keV.
The initial dataset contains 251 GRBs, including all classes. Compared with Dainotti et al. (2024c), there are an additional 30 GRBs with redshift that are added to the sample, an increase of 13%. From this sample, we remove 13 SGRBs and are left with 238 LGRBs with all the features listed above, as well as our response variable, the $\log(z+1)$ [^1] Note that all the logarithms in this work are in base 10. The response variable is the variable that we train our ML models to predict. We call this set of 238 GRBs the training set. This training set is 20% larger than the previous training set of Dainotti et al. (2024c). For the generalization set, we use the 221 GRBs of Dainotti et al. (2024c) and add 78 GRBs (an increase of 35%) taken from the Swift website. The additional 78 GRBs of the generalization set were fitted with the W07 model to obtain their $T_{a}$, $F_{a}$, $\alpha$, and $\beta$.
Figure 1: This figure depicts a typical GRB LC. The X-ray LCs are fitted with the W07 model, denoted with the darker red line. The end time of the plateau ($T_{a}$,$F_{a}$) and the prompt emission peak are shown as a black-filled dot and as an empty dot, respectively.
3.1 Comparison between the previous sample and the extended sample
In Fig.2 we present the scatter matrix plot between the training set from Dainotti et al. (2024c) (cyan) and the 30 new GRBs (orange) added in this analysis for comparing the distribution of the variables. We applied the Kolmogorov-Smirov test (KS-test) for each variable between the two data sets. All the variables pass the KS-test with a p-value $>$ 0.05, indicating that they originate from the same parent population. However, $T_{90}$ and $\gamma$ obtained p-values $<$ 0.05.
In the left panel of Fig. 3, the redshift distribution is illustrated, with the new sample depicted by a dashed line and the old sample by a solid line, both normalized to 1. The right panel of Fig. 3 presents the actual distribution. The redshift distributions pass the KS-test with a p-value of 0.35. Thus, we can say that the two sets of redshifts originate from the same parent population.
Figure 2: The scatter matrix plot shows the distribution of the features involved, with the diagonal showing the distributions. Here, the additional 30 GRBs added to the training set are orange dots, while the previous sample from Dainotti et al. (2024c) is represented by cyan dots. The numbers in the upper triangle show the correlation between two respective features, measured for new GRBs, the previous sample, and the combined sample.
Figure 3: Left panel: The redshift distribution of the new sample plotted in dashed and the old sample in continuous line normalized to 1. Right panel: the distribution of the new and the old samples not normalized.
4 Methodology and web app
In this research, we expand the methodology implemented in Dainotti et al. (2024c), and in addition, we make it available to the community via an easy-to-use and open-source web app. This web app is designed to let users input GRB parameters and quickly obtain redshift estimates within hours of the GRB trigger. This will allow the community to conduct follow-up observations of interesting bursts, such as high-$z$ events or low-$z$ low-luminosity GRB (LLGRB) events. This web app is also highly modular, and users are free to train their own custom models using their data.
We here detail in each subsection the various methods implemented in our analysis and the associated functions of the web-app.
4.1 MICE
Figure 4: The figure shows the distribution of missing data in our data set. The figure is divided into rows and columns. The top axis of the columns corresponds to the features in the data set, and the bottom axis corresponds to the total number of missing entries in that specific feature. Thus, for $\log(NH)$ there are 19 missing entries, for $\log(Peakflux)$ there are 5 missing entries, and so on. The left axis corresponds to the number of GRBs, and the right axis corresponds to the number of features with missing entries. The blue squares show the complete entries, and the red squares show which features are missing for a given number of GRBs. Thus, the first row indicates that there are 209 GRBs with no missing features, and the second row indicates that there are 15 GRBs with missing entries in only 1 feature, which is $\log(NH)$. The third row indicates that there are 3 GRBs with missing entries in 1 feature, which is $\log(Peakflux)$, and the fourth row indicates that there are 2 GRBs with missing entries in 2 features simultaneously, which are $\log(Peakflux)$ and $\log(NH)$.
Similar to the sample in Dainotti et al. (2024c), both our training and generalization sets have missing data. Out of the 238 LGRBs composing our training set, 29 GRBs have missing variables: 15 GRBs have missing $\log(NH)$ entries, 3 GRBs have missing $\log(Peakflux)$ entries, 2 GRBs have missing entries in both $\log(Peakflux)$ and $\log(NH)$ simultaneously, 3 GRBs have missing entries in $\alpha$, 2 GRBs have missing entries in $\alpha$ and $\log(NH)$ simultaneously, 2 GRBs have missing entries in $PhotonIndex$, 1 GRB has a missing entry in $\beta$ only, and 1 GRB has missing entries both $\gamma$ and $\beta$, simultaneously (see Fig. 4).
In Dainotti et al. (2024c), we removed the GRBs from the generalization set with missing entries. However, in this analysis, we impute them in a similar fashion to the training set. It should be noted that the imputation for the training set and the generalization set are conducted separately so that the training set does not influence the generalization set, thus preventing data bleeding. All 299 GRBs in our generalization set are LGRBs, out of which 11 have missing data entries. Two GRBs have missing entries in $\gamma$, two have missing entries in $\log(NH)$, two GRBs have missing entries in both $\gamma$ and $\log(NH)$ simultaneously, two GRBs have missing entries in $\log(Peakflux)$ and $\log(T_{90})$ and three GRBs have simultaneous missing data in $PhotonIndex$, $\log(T_{90})$ and $\log(Peakflux)$. In addition to the previously noted missing entries, we identify GRBs with $\log(NH)<20$, $\alpha$, $\beta$, and $\gamma>3$, and a $PhotonIndex<0$ as outliers, as these values fall towards the extreme ends of their respective feature distributions. Instead of removing these values from the analysis, we consider these as missing data and impute them alongside the previously mentioned missing data.
To minimize data loss and enhance the size of both the training and generalization sets, we utilized the imputation technique known as Multivariate Imputation by Chained Equations (MICE, Van Buuren & Groothuis-Oudshoorn (2011); Gibson et al. (2022)). MICE leverages the data currently in use to infer the missing data. The method applied within MICE is called Predictive Mean Matching (PMM, Little & Rubin (2019)). PMM initially fills the missing entries with the mean value of that particular feature. Here, we stress that random imputation is appropriate since the missing values of the variable are mainly due to the orbital gaps in the light curves. Then, it trains a regression model on the available data and, using this model, predicts the missing entries. Next, it finds the feature’s closest observed values to this prediction and randomly selects one of them as the imputed value. This process is carried out iteratively for each feature with missing entries until the values stabilize or a predetermined number of iterations is reached. We perform this iteration 20 times to have a reliable imputation. With the application of MICE, we increase our training sample by 14% and the generalization set by 4%.
We also implement MICE imputation in the webapp, so that any user can upload data with missing entries and obtain a complete data set with imputed values. The web app also shows the missing entries in the user’s data, as shown in Fig. 4 when the ’Apply MICE’ option is selected (see Fig. 8).
4.2 Machine Learning methods
In the upcoming sections, we outline this study’s machine learning methods and models.
4.2.1 LASSO feature selection
Figure 5: LASSO coefficients for all 10 predictors. The top 7 are chosen for further analysis.
The Least Absolute Shrinkage and Selection Operator (LASSO) is a method for selecting the best features for a given response variable. This method employs shrinkage for linear regression by constraining the $\ell^{1}$ norm (the sum of the magnitudes of all vectors in the given space) of the solution vector (in this case, $\log(z+1)$) to be less than or equal to a positive value called the tuning parameter ($\lambda$). This penalization enables the model to choose a subset of features and eliminate the remaining by assigning a zero coefficient (Tibshirani, 1996). In this work, we apply the LASSO selection feature using the GLMNET function (Hastie et al., 2017; Tibshirani et al., 2012; Dainotti et al., 2021a). Then, we select the features’ coefficients corresponding to the largest value of $\lambda$, with errors less than 1 standard deviation (Friedman et al., 2010). The top seven features, $\log(NH),\log(Peakflux),PhotonIndex,\log(T_{a}),\log(F_{a}),\gamma$, and $\alpha$, (see Fig.5) are picked for further analysis.
4.2.2 The ML Models used
We tested multiple different ML models to find the best set of models that can give accurate redshift predictions. Based on our analysis, we determined three models that perform best, which are described below.
  • Generalized Additive Model (GAM, Hastie & Tibshirani (1990)) is a semi-parametric method in which the response variable is connected to the predictors via the sum of either a parametric or non-parametric function or a combination of both. GAM provides the option for fitting smooth spline functions to the data that can enable non-linear fits. We explore the performance of GAM with and without the application of smoothing functions.
  • Generalized Linear Model (GLM, Nelder & Wedderburn (1972)) is an extension of the linear model that uses an iteratively weighted least squares method for the estimation of maximum likelihood to fit the relation between the response variable and the predictors. GLM allows for non-linear distributions of the response variable through a link function. In our analysis, we use the default Gaussian link function.
  • Random forest (RF) is tree based ML method that generates multiple independent regression trees (Breiman, 2001; Miller et al., 2015; Dainotti et al., 2021a). The final results are an average of these regression trees. RF has three hyperparameters: depth, number of trials, and number of trees. Hyperparameters are function parameters that determine a function’s properties and dictate how well they fit a given set of data. In our analysis, we use the default implementation of RF from the Caret package (Kuhn & Max, 2008).
4.2.3 Formula Generation
Following the methodology established by Dainotti et al. (2024c), we aim to select the best formulas for GAM and GLM that can give us the most accurate predictions.
After performing LASSO-based feature selection to determine the set of the most predictive variables, we apply a method we call the formula generator. This method creates formulas based on combinations of these seven predictors and their squared terms. The sets of best formulas for GAM and GLM are determined independently. The process is described below. One of the formulas that perform best with GLM is shown in Eq. 1 for demonstration purposes. The full list of formulas used in this analysis can be found on the web app and the text files provided in the associated GitHub repository.
To determine the best formulas for our data set, we divide the set of GRBs with $z_{obs}$ into a training and test set with an 80:20 split. A total of 16510 formulas were created from combinations of the seven variables chosen by LASSO (see Sec. 4.2.1). Each formula was tested by performing a 10-fold cross-validation (10fCV) with the training set 100 times. 10fCV is an ML method for assessing the performance of a model. It splits a given training set into 10 parts (folds). The model is trained on 9 folds and predicts on the 10th fold. This step is repeated until the model has predicted for all 10 folds, and then the results are assimilated, providing a measure of the model’s performance. We thus obtain cross-validated Pearson correlation coefficient ($r$) and root-mean-square (RMSE) values for each formula. Based on these scores, we select a subset of formulas that have achieved $r$ above 99.9% quantile and RMSE score below 2% quantile (see red points in the left panel of Fig. 6). The exact cutoff for $r$ is 0.553, and for RMSE is 1.145. This subset of formulas is used for obtaining predictions on the test set, which we had initially created (see right panel of Fig. 6). Finally, we select the best formulas based on this test set’s performance. Specifically, we pick three formulas that achieve the highest $r$, lowest median absolute deviation (MAD), and lowest RMSE.
Figure 6: Left panel: Plot showing the RMSE and Correlation scores for 16510 formulas. The red points show the formulas selected for analysis on the test set. These formulas achieve correlation and RMSE scores above the cut-offs mentioned at the top of the figure. Right panel: Performance of the selected subset of formulas on the test set. Each plus symbol here represents one formula selected from the left panel (red points). The red vertical and horizontal lines point to the formula that obtained the lowest RMSE and highest correlation on the test set.
[…]
However, during our analysis, we discovered that the formulas determined via the above-mentioned methodology are sensitive to the particular training set selected. Thus, minimizing this strong dependence on the training set and selecting formulas that generalize to the full set of GRBs is important. To accomplish this, we follow the steps mentioned above; however, we randomize the GRBs used in the training and test sets while maintaining the 80:20 split ratio. We created 100 pairs of training and test sets and performed the subsequent steps for each pair. Furthermore, all the above-mentioned steps were repeated independently for GAM and GLM.
At the end of the computation, we counted which formulas were the best on the 100 different test sets and selected those that appeared the maximum number of times. With this extension to the methodology, the formulas picked in this manner are generalized to the full GRB data set rather than a particular instance of training-test set combination. Following this, we pick six formulas for GAM and four formulas for GLM.
In addition to testing the 16510 formulas with GAM and GLM, we also tested exclusively 148 formulas with GAM that have smoothing terms. We note here that even though these 148 formulas with smoothing were included in the 16510 formula set, they were not chosen during the analysis. However, it is essential to test these because in our previous work, (Dainotti et al., 2024c), smoothing formulas worked better, as they can capture the nonlinear relations between features. Thus, we decided to test them separately. These 148 formulas include smoothing terms applied to features individually and in pairs. The methodology described above was applied to these 148 formulas, and we obtained two formulas that work best with our data set. These were included in our final ensemble.
We also tested formulas where three and four features were applied to the smooth function together. However, this did not improve our ML performance, and thus, we did not go forward with that set of formulas.
On the web app, we enable users to obtain the best formulas for their given data set. However, the full methodology mentioned above has not been implemented due to computational limitations. Performing 10fCV on 100 randomized training test sets for thousands of formulas requires extensive computational time. We leveraged the supercomputing cluster at the Center for Computational Astrophysics at NAOJ, Tokyo, Japan and the PLGrid facility at Cyfronet, Krakow, Poland for our analysis and results. This, however, is beyond the scope of a web application at this stage. The formulas provided by the web app are determined on a single randomized training test set, similar to Dainotti et al. (2024c). Furthermore, the formulas are generated using only the variables, not their squared terms. However, formulas with smoothing terms are included within this.
The web app, however, allows for the customization of correlation and RMSE thresholds to determine the formulas that meet the user’s criteria. However, as this work demonstrated, it is advisable to select a higher percentile for correlation and a lower percentile for RMSE. The idea behind the chosen percentile is that we would need to keep the number of selected formulas small to prevent over-fitting. The web app determines GAM formulas by default. However, the user can choose GLM by selecting the ’Use GLM’ option. The interface for the web app is presented in the upper panel of Fig. 7 and the final formulas are presented in a tabular format, as presented in the bottom panel of Fig. 7.
Refer to caption Refer to caption
Figure 7: The figure shows the GAM/GLM Formula Generator page of the web app. This module allows users to access the best formula for predicting redshift based on the training set provided and its predictors. The user can also manually set the cut-off for the correlation and RMSE for selecting the formula.
4.3 Outlier removal
Before performing any ML training, we need to remove outliers, as these can affect the model performance. We use the M-estimator technique to remove outliers (Huber, 1996, 1964). This is a robust regression method where the outliers are determined based on a robust loss function. Unlike ordinary least squares regression, which is sensitive to outliers, M-estimator assigns lower weights to extreme observations, making the results more robust. For the outlier removal, we selected the formula (Eq. 1), which obtained the highest counts for GLM. Using this formula, the M-estimator fits the 238 GRBs and assigns weights ranging from 0 to 1. GRBs that obtained weights less than 0.65 were eliminated. This specific weight of 0.65 was selected because it removes 5% of the data set as outliers. Thus, we are left with 226 GRBs, which are split again into training and test sets containing 180 and 46 GRBs, respectively. This training set of 180 GRBs is 20% larger than the training set used in Dainotti et al. (2024c).
For the generalization set, we remove GRBs that lie outside the training set’s parameter space. This minimizes the ML model’s chance of extrapolating results, which may lead to inaccurate redshift predictions. Thus, using our trained model, we predict the redshifts of this trimmed generalization sample of 276 GRBs.
4.4 The Superlearner
SuperLearner is an ensemble method that combines multiple ML models and leverages these models to obtain the best prediction. SuperLearner assigns to each model a weight based on its performance, and the sum of all weights is 1. The performance of the constituent models is determined via the 10fCV method, which is built into SuperLearner. The advantage of the SuperLearner is that its prediction is a weighted average of the predictions of its constituent models. We also use SuperLearner weights to select models that perform the best and models that obtain higher weights are included in the final nested 100 times 10fCV procedure. This procedure provides us with the results presented in this paper.
In our web app, we allow users to provide their data and obtain results of the SuperLearner through the metrics of our analysis: the Pearson correlation plots, both in linear and log scale, between $z_{obs}$ and $z_{pred}$, the bias, defined as $<z_{pred}-z_{obs}>$, the normalized median absolute deviation (NMAD), and the RMSE. Additionally, the SuperLearner module of the web app also supports features such as M-estimator, MICE, the removal of catastrophic outliers, model integration such as GAM, GLM, etc., and parallel processing. This enables users to tune the functionality to their liking. The user interface of the Superlearner module in the web app is presented in Fig. 8.
Figure 8: The figure depicts the SuperLearner page of the web app, showing the data upload panel and options to choose model and data parameters.
The results presented here use a SuperLearner model that incorporates six GAM formulas, two smooth GAM formulas, four GLM formulas, and RF (see Sec. 4.2.2 and Sec. 4.2.3 for details).
4.5 Balance sampling
We aim to investigate how a more balanced sample of sources will affect the performance of our ML model. A balanced sample is a data set in which the response variable’s distribution is close to uniform. Such a distribution helps train the model over the full range of potential values for the response variable, thus helping it generalize to future samples. In this analysis, we deal with an unbalanced sample where the number of GRBs at $z>2.5$ (80) is fewer than the number of GRBs at $z<2.5$ (158). This phenomenon can be attributed to the Malmquist bias effect, which causes us to only detect the brightest astrophysical sources at high-$z$ due to detector thresholds. This effect prevents us from seeing a large population of dim sources at high-$z$. If we had a perfect satellite with extremely high sensitivity, we could populate the distribution at high-$z$. There are several methods to balance the sample: up-sampling, down-sampling, and balance sampling. For up-sampling, the original data remains unchanged, and duplicate samples of the minority class are added. In down-sampling, data is randomly selected to ensure that all classes are equally represented, matching the frequency of the minority class.
For up-sampling, multiple methods exist. We applied the Synthetic Minority Over-sampling Technique (SMOTE, Chawla et al. (2011); Dal Pozzolo et al. (2013)). SMOTE, as the name implies, creates synthetic copies of the minority class, thereby increasing its sample count. This is achieved by generating new data points within the feature space of the minority class dictated by the real nearest-neighbor data points. This prevents overfitting, which can happen when minority class samples are duplicated. It also eliminates the need to remove samples from the majority class, which may lead to information loss and reduced accuracy.
The redshift distribution of the upsampled data is shown in Fig. 9. This balanced sample also undergoes the same data treatment process as mentioned in Sec. 4.3.
Since this is an important part of the inference of the redshifts, we have also made this part available through the web app, see Fig. 8. The user can tick the ’Apply Upsampling’ button to obtain results with an augmented and balanced sample.
Figure 9: The redshift distribution obtained with the upsampling and the original redshift distribution are shown in solid and dashed lines, respectively.
4.6 Bias correction
Bias correction is a technique that is used to correct biases in an ML model’s prediction.
Bias is defined as the difference between the observed and predicted response variables. ML models trained on an imbalanced training set are more biased towards the range containing more examples. In our sample of GRBs, the models tend to predict more toward lower redshift ranges as there are more GRBs in that range (see Sec. 4.5). Balanced sampling can mitigate this issue to an extent. However, we apply the Optimal Transport bias correction technique (Ferradans et al., 2013) to address this issue more directly. Some of us have already used it in Dainotti et al. (2024c, b); Narendra et al. (2022).
In this technique, a linear model is fitted to the sorted values of observed and predicted $\log(z+1)$. The predicted values are corrected by multiplying the slope and adding the intercept of the linear fit. We perform this bias correction on the 10fCV results, and the generalization set predictions.
4.7 Monte Carlo analysis
With the MC simulations, we aim to explore the reason for the changes in our results compared to Dainotti et al. (2024c). The MC simulations also help us assess the degree of uncertainties in the prediction once the sample size is increased. Since in the procedure detailed in Dainotti et al. (2024c), we have used MICE and then run the Superlearner on the imputed data, we here perform MC simulations on the imputed data to keep the comparison more straightforward.
We perform simulations on the 7 predictors which are selected by LASSO, see Fig. 5. We simulate three sample sizes: 14, 28, and 42. This reflects, on average, how many GRBs with redshift and plateau will be added to the Swift catalog in one, two, and three years, respectively. This average was calculated based on the training set currently being used. Each sample size was generated 200 times with unique instances and combined with the training set of observed GRBs. For sampling, we used the SMOTE method (see Sec. 4.5), as this enables multivariate sampling of the data. Multivariate sampling is important as it maintains the correlations between the seven parameters involved, such as between $\log(T_{a})$ and $\log(F_{a})$ and $\log(Peakflux)$ (Dainotti et al., 2008). We simulated the samples so that the distributions of the variables resemble the current ones, thus we retain only the samples that pass the KS test. Each simulated sample was tested with 10fCV using the ensemble SuperLearner model described in Sec. 4.4. The results from this analysis are presented in Sec. 5.4.
4.8 Model Comparison with web-app
The web app’s Model comparison module is designed to allow users to investigate their data using more ML models than those used in this work. This is important because the best model for predicting redshift may vary depending on the data being used. Figure 10 shows the Model Comparison module’s interface on the web app. This process is initiated by importing data files provided by the user. Moreover, this module also offers functionalities such as MICE for data imputation and allows users to specify the number of loops for 10fCV (see Sec. 4.2.3). The web app lists predefined models for use, as shown in Fig. 10. Users can tailor their selections based on specific requirements and commence the analysis by clicking on the submit button. The application then constructs an ensemble SuperLearner model based on the selected parameters and models. The application generates a correlation plot depicting the $z_{pred}$ vs $z_{spec}$ distribution and metrics showing the performance of the selected models over the provided dataset. This visualization aids users in understanding the relative strengths and weaknesses of the models under consideration.
Figure 10: This figure shows the various options and parameters available to a user of the web app when using the Model Comparision functionality.
4.9 Relative Importance of variables
The Relative Importance module of the web app evaluates the relative significance of the variables used for predicting the redshift. The process is initiated by importing data files the user provides. The ML model used for this module is the best model determined from our analysis in this publication (see Sec. 4.2.2). The relative importance of the predictors is decided based on this model’s performance. Users can also use the pre-trained models from the web app’s Model Comparison or Superlearner modules to obtain relative importance values for their data.
The interface of the web app for this module is presented in the upper panel of Fig. 11.
Refer to caption Refer to caption
Figure 11: Upper panel: The user interface of the Relative Influence module of the web app. Lower panel: Relative importance bar plot for the predictors used in this analysis.
4.10 Data Visualization with the web app
The Data Visualization module of the web app provides the user with an option to visualize their training set or generalization set via the scatter matrix plot (as shown in Fig. 2) and missing data (as shown in Fig. 4). The scatter matrix plot offers a comprehensive view of the uploaded data along with histograms and correlation coefficients between pairs of features. This visualization tool facilitates a deeper understanding of the dataset’s relationships and distributions, aiding users in making informed decisions and gaining valuable insights. This module’s plot of missing data helps users determine which features in the provided data set have missing entries.
Figure 12: This figure shows the interface for the Data visualization module on the web app.
5 Results
Our Results section is split into four subsections. First, we present the results demonstrating the effectiveness of our ML model. Here, we also show the results obtained using balanced sampling. Second, using the trained model, we present the redshift predictions obtained on the generalization set. Third, we present the results from our MC analysis. Finally, we provide an overview of the performance of the web app’s redshift estimator module.
5.1 Performance of the ML model
The results shown here are obtained after performing 10fCV 100 times on 180 GRBs using the SuperLearner model as described in Sec. 4.4. The correlation plot between $z_{obs}$ and $z_{pred}$ obtained from SuperLearner is shown in the upper left panel of Fig. 13. We achieve an $r$=0.646, RMSE=1.011, bias=0.14 and NMAD = 1.34. We apply the optimal transport bias correction (see Sec. 4.6) to these results to obtain the plot shown in the upper right panel of Fig. 13. Here, we obtain $r$=0.89, RMSE = 0.62, bias = 0.0047, and NMAD = 0.86. Sec. 6 discusses how these results compare with those presented in Dainotti et al. (2024c) and its implications.
We also present the results for our ML model when using a balanced sample. The plots are presented in the bottom left and right panels of Fig. 13. The bottom left panel shows the correlation plot obtained from the SuperLearner model, similar to the upper left panel without balanced sampling. We achieve $r$ = 0.668, RMSE = 1.005, bias = 0.13, and NMAD = 1.56. The results then undergo the same bias correction step, and they are shown in the bottom right panel of Fig.13. Here we obtain an $r$= 0.902, RMSE= 0.59, bias= 0.003, and NMAD= 0.76. The discussion about these results compared to those without balanced sampling is presented in Sec. 6.
Ukwatta et al. (2016) obtained $r=$0.57 using RF. We have the four features common with their analysis, namely $T_{90}$, $PhotonIndex$, $\log(Fluence)$, and $\log(NH)$. Our correlation value is 13.3% higher for the SuperLearner results and 56% higher for the bias-corrected results. Our balanced sample results are 17.1% higher, and the bias-corrected results are 58% higher. We can attribute this improvement to our use of a more robust methodology and the plateau features. We also obtained similar results in Dainotti et al. (2024c).
Aldowma & Razzaque (2024) use DeepNeural networks and RF with Fermi-GBM and Konus-Wind observed GRBs for estimating pseudo-redshifts. They achieved the highest $r^{2}$ of 0.86 on their test set, which is equal to $r=$0.92. Compared to our bias-corrected results, they achieved an improvement of 4%. However, in their analysis, it is unclear whether they discard duplicated GRBs between the Konus-Wind and Fermi-GBM catalogs. Furthermore, they seem to have trained the model on a single training-test split. These two effects can make the ML model overfit their data, leading to a higher correlation score.
Racz (2023) used RF and extreme gradient boosting with GRBs taken directly from the Swift catalog. Similar to our analysis, they use $\log(z+1)$ as their response variable and obtain a $r=$0.76 in the $\log(z+1)$. However, other metrics such as RMSE, bias, and NMAD are not quoted. Furthermore, they do not quote the metrics in the linear scale. Thus, comparing our results in the $\log(z+1)$ scale, we find that our $r$ is lower by 20% and higher by 19% for the non-bias-corrected and bias-corrected results, respectively.
Figure 13: Results obtained from performing 100 times 10fCV using the SuperLearner model. The bottom and upper panels present results with and without balanced sampling, respectively. The upper left panel shows the result obtained from the SuperLearner model. The upper right panel shows the corresponding bias-corrected results. The bottom left panel shows the result obtained from the SuperLearner model when using a balanced sample. The bottom right panel shows the corresponding bias-corrected results.
5.2 Relative importance results
The relative importance of the seven features is presented in the lower panel of Fig. 11. Similar to Dainotti et al. (2024c, b), we observe that the highest influence on the redshift prediction is assigned to $\log(NH)$. The $\log(NH)$ is obtained from the Swift catalog and is determined by spectral fitting with a fixed galactic absorption with the HEASoft software [^2] https://heasarc.gsfc.nasa.gov/ftools. Thus, the $\log(NH)$ value we use here is the column density between the burst and the Milky Way.
The dependence of $\log(NH)$ on the redshift is expected; however, this relation is non-trivial. The value of $\log(NH)$ is impacted by the density and metallicity of the intervening medium, and thus, we discuss some of the factors that can affect this complex correlation(Arumaningtyas et al., 2024).
  • The first factor is the absorption of the GRB’s emission by the intergalactic medium between our galaxy and the GRB’s host galaxy. Evidence for this was found by Rahin & Behar (2019), who attributed the evolution of $\log(NH)$ with redshift to the accumulation of the ionized and diffuse gas in the intergalactic medium.
  • The second factor is the interstellar medium (ISM) density inside the host galaxy. Generally, high-$z$ galaxies have larger star-forming regions, and this corresponds to the presence of more neutral hydrogen in the ISM. This causes more absorption of the GRB’s emission than inside low-$z$ galaxies, thus affecting the $\log(NH)-z$ relation.
  • The third factor is the evolution of the metallicity of the progenitor’s ejecta. Heintz et al. (2023) found that the effective metallicity derived from the GRB spectra does evolve with redshift, and thus, it might have some impact on the $\log(NH)-z$ relation. We expect the progenitors of GRBs at high-$z$ to have lower metallicity than their low-$z$ counterparts. This can lead to the complex evolution of absorption of the GRB’s emission and, thus, significantly affect the $\log(NH)-z$ relation.
  • Finally, the metallicity of the host galaxy’s ISM has a limited impact on the $\log(NH)-z$ relation. This is evident from the findings of (Cucchiara et al., 2015; Graham et al., 2023) where they observed that high-$z$ GRBs occur in low metallicity regions. Furthermore, Bolmer et al. (2018) found that GRB host galaxies at $z>4$ have less dust and extinction value ($Av<0.5$mag) than low-$z$ sources. This leads to the conclusion that the evolution of metallicity cannot be explained by higher dust content at $z>4$.
Going back to the relative importance, then the second most predictive feature here, and in Dainotti et al. (2024c), is $\log(T_{a})$. The plateau features are derived using the X-ray LCs in both cases. This differs from the second most predictive variable found in Dainotti et al. (2024b), which was $\log(F_{a})$. The plateau features were obtained from optical LCs there. The fact that $\log(T_{a})$ is correlated to $\log(F_{a})$ implies that the difference between the X-rays and optical results is indeed not significant. The discussion about the implication of these results is presented in Sec. 6.
5.3 Estimating new redshifts in the generalization sample
As stated in Sec. 3, we have increased the generalization sample by 35%, from 221 GRBs (Dainotti et al., 2024c) to 276 GRBs. The histogram plots comparing the training and generalization sets’ features are shown in Fig. 14. We compared the distribution of each parameter between the generalization set and the training set via the KS test, and the p-values are stated at the top of each parameter distribution.
Figure 14: Histograms showing the distribution of the seven features of the generalization set against the training set. The KS test p-value scores are shown in the title of each figure. The dashed lines and solid lines represent the generalization and training sets, respectively.
Compared to the training set, all distributions, except the one of $\log(NH)$, have a p-value greater than 0.05. This means that the features are sampled from the same parent distribution. This indicates that the training set’s distribution of $\log(NH)$ differs from that of the generalization set and is not representative of the parent population.
Figure 15: This figure shows the scatter matrix plot of the generalization and training sets. The generalization set is presented with orange dots, and the training set is presented with cyan dots. The diagonal shows the histogram distribution for the generalization set, while the upper triangle shows the Pearson correlation values for respective pairs of parameters.
In Fig. 16, we compare the distributions of predicted redshift and the redshift of the training set. The training set’s redshift distribution is presented in red histograms, and the generalization set’s redshift distributions are presented in white and green histograms, with and without bias correction, respectively. Performing the KS test between the redshift distributions of the training set and the generalization set gives p-values lower than the threshold of 0.05. Specifically, we obtain p-values of $2.5\times 10^{-7}$ and $0.0002$ for the non-bias-corrected and bias-corrected redshift distributions, respectively. This discrepancy can be attributed to the different $\log(NH)$ distributions of the generalization and training sets. However, we can observe that the bias correction step does bring the redshift distribution closer to the training set distribution. Thus, this is an essential step in the redshift prediction.
It should also be noted that GRBs with and without $z_{obs}$ might differ in their physical origin and chances of observations. Thus, they can have different intrinsic redshift distributions. The low-luminosity sources have a lower chance of having their redshift measured due to the difficulty of obtaining the redshift from the afterglow spectral features, the trouble of observing the host galaxy, or the fact that the host galaxy is not in the database and is not pointed by any observing facilities. Moreover, we observe a statistically smaller $\log(NH)$ in the generalization set. This corresponds to less absorption via the intergalactic medium, thus making the absorption-based redshift measurements harder. Those two effects could explain the difference in redshift distributions between the training and generalization sets.
We show a boxplot distribution of the predicted redshifts in the lower panel of Fig. 5.3. The red and grey boxplots show the bias-corrected and non-bias-corrected distributions. The x-axis shows the names of the generalization set GRBs, and the y-axis shows the redshift values. Note that because of the image’s scale, not all the GRB names are present along the axis. Our work here has enabled us to increase the Swift sample with redshift estimates by 110%.
Figure 16: Upper panel: The redshift distribution of the generalization and training sets. The training set is shown with red histograms, while the generalization set with and without bias correction is shown in white and green histograms, respectively. Lower panel: This figure shows the box plot distribution of the generalization set redshift predictions, sorted from lowest to highest redshift. The bias-corrected and non-bias-corrected distributions are shown in red and grey boxplots, respectively. The dots show the outlier predictions obtained.
5.4 Monte Carlo simulation results
To understand if the result of an MC simulation is due to one realization only, we have then repeated the simulations with 200 different samples to de-randomize the process. The MC simulation results shown in Fig. 17 combine 200 instances of 14, 28, and 42 samples. Following the 10fCV analysis of each sample, we plot the $r$, RMSE, bias, and NMAD distributions for each sample size in columns 1, 2, and 3 for the samples of 14, 28, and 42, respectively. The red vertical line in the $r$ and RMSE distributions shows the mean of the distribution.
Figure 17: This figure shows the results of our MC analysis. All the plots shown are combined results of 200 simulated samples. The first column of histograms shows the distribution of properties for the 14 sample cases. The second and third columns of histograms present the distribution of properties for 28 and 42 sample cases, respectively. The first row shows the correlation distribution, and the red line shows the mean. The second row shows the RMSE distribution. The third and fourth rows show the bias and NMAD distributions, respectively.
The 14-sample case: The top image shows the $r$ distribution. Here, we see the mean value of the distribution is 1.5% higher at $r$=0.656 than the $r$ value of our SuperLearner results (see the upper left plot of Fig. 13). In fact, 80% of the simulated samples attained a $r$ greater than our SuperLearner result of $r$=0.646. The mean value of RMSE is 3% lower at RMSE=0.978 than our SuperLearner result.
The 28-sample case: The mean values of the $r$ and RMSE distributions are 2.3% higher at $r$=0.661 and 3.7% lower at RMSE=0.961, respectively, than our SuperLearner results. Here, 94% of the simulated samples attained a $r$ greater than our SuperLearner result.
The 42-sample case: Here, we see that the mean values of the $r$ and RMSE distributions are 3.4% higher at $r$=0.668 and 4.4% lower at RMSE=0.956, respectively, than our SuperLearner results. Also, 97.5% of the simulated samples attained a $r$ greater than our SuperLearner result.
Based on the above-mentioned results, we can say that for the majority of simulated samples, there is an improvement in the results across all metrics. This result is very promising because it indicates that our ML model generalizes to the GRB sample well and should give improved results for at least three years.
5.4.1 Future of Multi-wavelength Redshift Estimates via Machine Learning
The following subsection draws upon the results from Sec. 5.4 and Sec. 5.3 for estimates on the number of GRB redshifts that can be predicted using ML analysis for future X-ray and optical observations of afterglows. These estimates are based on the estimates presented in (Dainotti et al., 2022b), which in turn rely on the existing X-ray and optical samples determined by Dainotti et al. (2024c) and Dainotti et al. (2024b), provided up to May 2025.
Table 1: The table details the number of GRBs this study will enable to have redshift predictions in the future using Swift observations.
For the X-ray afterglow, we currently have 255 GRBs possessing plateaus and measured $z_{obs}$, observed from 03/2005 to 12/2023. We have 299 events with X-ray plateaus but without redshifts, observed from 01/2005 to 02/2024. Both sets of GRBs are taken from the Swift/XRT catalogs. Based on these data, $\sim$13.6 GRBs with X-ray plateaus and $z_{obs}$, and $\sim$15.7 GRBs with X-ray plateaus and unknown redshifts are expected annually. For GRBs with $z_{obs}$ and X-ray plateau, 24 additional GRBs are expected from 12/2023 until 09/2025. For GRBs with X-ray plateaus and without $z_{obs}$, 25 additional GRBs are also expected from 02/2024 until 09/2025. This sample with unknown redshift will be the generalization set. So, the number of GRBs with $z_{pred}$ from ML analysis will be $>$324.
For the optical afterglow,e have gathered 536 LCs through the 416 ground-based Telescopes, the HST, the Swift/UVOT satellites, the GCN, and private communications. These 536 GRBs, observed from 05/1997 to 10/2023, have $z_{obs}$ and optical afterglow. Out of these 536 GRBs, 180 (34%) GRBs have optical plateau observations. As of this writing, according to the Greiner webpage [^3] https://www.mpe.mpg.de/ jcg/grbgen.html, the total number of GRBs with optical afterglow is 932, and 622 have secure redshifts. Thus, 310 GRBs (932 - 622) do not have redshift but have an optical afterglow. Assuming the same fraction of plateau observations in the optical afterglow, our current generalization set would be 34% of 310, which is 104 GRBs. If we extrapolate to 09/2025, we expect to have 112 GRBs with an optical plateau and no measured redshifts. Similarly, we expect to have 193 GRBs by 09/2025, which have an optical plateau and $z_{obs}$. Thus, we expect $>$112 GRBs with $z_{pred}$ from ML analysis. Thus, the new $z_{pred}$ will be $>$436 (324+112), where 324 and 112 are GRBs with X-ray and optical plateaus, respectively. With this increase of 92% in the redshift sample, we will enhance the accuracy of the LF and density rate evolution estimates, thus helping to address the discrepancy between the GRBFR and SFR evolution. The above numbers are also reflected in the Table 1. This analysis can be extended for future reference, and it is our basis to compute how many GRBs we will have approximately in the next one, two, and three years from now.
Dainotti et al. (2022) also calculated the number of GRB observations required to achieve similar precision on $\Omega_{M}$ for the flat $\Lambda$CDM model as obtained in studies by Conley et al. (2011); Betoule et al. (2014); Scolnic et al. (2018) (for SNe Ia).
According to Table.9 of Dainotti et al. (2022), to reach the precision of Conley et al. (2011), the lowest number of X-ray platinum sample GRBs with redshift needed, with half error bars on plateau parameters, is 354. To achieve this number, we assume that by using an updated lightcurve reconstruction (LCR) technique, an improvement of the methodology from Dainotti et al. (2023), which is in preparation, we will have the same number of GRBs with half error bars (denoted by n=2) as with full error bars (denoted by n=1). Currently, we have 105 GRBs that belong to the platinum sample with $z_{obs}$. Assuming the same fraction of platinum GRBs exists in the generalization set, we should have another 125 GRBs from the generalization set used in this work. Combining these two, we estimate that we currently have 230 platinum sample GRBs with redshift estimates. Further, considering the observations we will obtain from SVOM (Godet et al. (2014) and SVOM webpage [^4] https://www.eoportal.org/satellite-missions/svom-eclairs-gamma-ray-imager) and the application of ML and LCR, we can expect to reach the precision of Conley et al. (2011) by 2029 (see PLATtrim20 column of Table.2). Thus, this ML effort reduces the number of years required to reach the precision of Conley et al. (2011) by 26 years. Similarly, to reach the precision of Betoule et al. (2014), we need the observational capabilities of THESEUS. According to our estimates in Table.2, again, using LCR and ML to boost the numbers and GRBs with half error, we should be able to reach the precision of Betoule et al. (2014) by 2041 at least.
Similarly, let us consider the optical estimates from Table 9 of Dainotti et al. (2022). We should be able to reach the precision of Conley et al. (2011) with the current data we have if we assume that we will have an equal number of half error GRBs as we have full error GRBs, achieved by using updated LCR methodology. If we, however, trim the sample in the closest 25 to the 3D fundamental plane (OPTtrim25 column of Table.3), we would have reached the precision of Conley et al. (2011) even in 2005, 6 years earlier than the one reached by the SNe Ia. Note that the numbers mentioned along the fourth row of Table.3 are estimated based on GRBs observed from January 2005. This is done to avoid negative numbers, as some of the estimates are reached well before 2024. The corresponding year estimates are obtained based on the rate of observation using Swift and other ground-based optical observatories. Furthermore, if we take into account the rate of observation of SVOM and the fraction of optical GRBs with the plateau we currently have, then we can reach the precision of Conley et al. (2011) by 2028 (only 4 years from now), with the current errorbars. However, to reach the precision of Betoule et al. (2014), we need to consider the observation of THESEUS as well. Considering THESEUS’s rate of observation to be 700 GRBs per year (Dainotti et al., 2022; Amati et al., 2018a), we again consider the same fraction of optical plateau GRBs as we currently have. With this, we can reach the precision of Betoule et al. (2014) by 2039, 7 years after the launch of THESEUS (the current expected launch is 2032, if the mission is approved). However, to reach the same precision as (Scolnic et al., 2018), we would need to wait at least until 2041 and use both ML and LCR and half error bars.
However, we can say that these estimates are rather conservative since new optical telescopes and facilities will be operating in the future, and we can rely on the efforts of some of us to continue building the largest optical GRB catalog to date (Dainotti et al., 2024a). Indeed, one of the outcomes of gathering multiple data points from several sources has the advantage of allowing an increase of data points, in general, in the LCs and, in particular, in the plateau region when it occurs. Thus, we assume that once the optical GRB sample is updated, our estimates will be larger in terms of the total number of optical plateaus, and thus, the number of years needed to reach SNe Ia precision will be less. In addition, regarding the X-ray sample, some of us are currently working on choosing a subsample, not driven by phenomenological reasons but driven by the magnetar model (Zhang & Mészáros, 2001; Rowlinson et al., 2014b; Yi et al., 2017). In such a case, we can have a smaller scatter of the plane as shown for the 2D Dainotti relation between the X-ray luminosity at the end of the plateau and its rest-frame duration Dainotti et al. (2008, 2011); Dainotti213; ainotti2017; Wang et al. (2022), and we can reach again the precision of the SNe Ia sooner, using the X-ray sample.
Table 2: This table presents the number of platinum sample Xray GRBs with redshift needed to reach the precision of $\Omega_{M}$ as obtained in studies by Conley et al. (2011); Betoule et al. (2014); Scolnic et al. (2018) (for SNe Ia). This is an extension of the table provided in Dainotti et al. (2022). n=1 indicates that the error bars on the GRB plateau parameters are full-width, while n=2 indicates that the error bars are halved. LCR indicates that the estimates for the number of GRBs have been boosted by using LCR, while ML indicates that the estimates have been boosted by using machine learning.
Sample PLAT PLATtrim10 PLATtrim20
Reaching Precision of Conley (before launch of THESEUS)
Original estimates (n=1) 789 847 646
n=2 + LCR + ML 127 169 124
Years to reach (n=2 + LCR + ML) 2029 2031 2029
Reaching Precision of Betoule (after launch of THESEUS)
Original estimates (n=2) 1452 1788 1466
n=2 + LCR + ML 1040 1376 1054
Years to reach (n=2 + LCR + ML) 2041 2044 2041
Table 3: This table presents the number of optical GRBs with plateau and redshift needed to reach the precision of $\Omega_{M}$ as obtained in studies by Conley et al. (2011); Betoule et al. (2014); Scolnic et al. (2018) (for SNe Ia). This is an extension of the table provided in Dainotti et al. (2022). n=1 indicates that the error bars on the GRB plateau parameters are full-width, while n=2 indicates that the error bars are halved. LCR indicates that the estimates for the number of GRBs have been boosted by using LCR, while ML indicates that the estimates have been boosted by using machine learning.
Sample OPT OPTtrim10 OPTtrim25
Reaching Precision of Conley (before launch of THESEUS)
Original estimates (n=1) 271 330 244
n=1 + LCR + ML 99 158 72
Years to reach (n=1 + LCR + ML) 2032 2037 2030
Original estimates (n=2) 142 112 36
n=2 + LCR* 30 79 3
Years to reach (n=2 + LCR) 2028 2024 2005
Reaching Precision of Betoule (after launch of THESEUS)
Original estimates (n=1) 1031 829 685
n=1 + LCR + ML 770 568 424
Years to reach (n=1 + LCR + ML) 2044 2041 2039
Original estimates (n=2) 284 393 350
n=2 + LCR 118 227 184
Years to reach (n=2 + LCR) 2035 2037 2036
n=2 + LCR + ML 23 132 89
Years to reach (n=2 + LCR + ML) 2032 2034 2033
Reaching Scolnic Precision (after launch of THESEUS)
Original estimates (n=2) 1086 1513 822
n=2 + LCR + ML 825 1252 561
Years to reach (n=2 + LCR + ML) 2045 2052 2041
5.5 Applicability to future missions
It is important to note that all the methodologies spelled out here can be applied not only to GRBs observed by Swift, but also to any GRB for which we can observe the plateau emission. Thus, we envision this method can be applied to GRBs observed by Space-based multi-band astronomical Variable Objects Monitor (SVOM,Wei et al. (2016)), the Transient High-Energy Sky and Early Universe Surveyor (THESEUS) mission (Amati et al., 2018b), and the Einstein Probe mission(Yuan et al., 2015). Surely, with the new data from other satellites, one needs to be cautious about how the variables’ parameter spaces will change compared to the space of the Swift variables and how much this will impact the prediction. Properly leveraging the power of these new satellites will require a re-training of the ML models. However, the base methodology described here and in our previous works (Dainotti et al., 2024c, b) should still produce reliable results.
The SVOM mission, in particular, is very promising indeed, as one of their key aims is to quickly identify and observe the afterglow of GRBs at both X-ray and visible wavelength. The SVOM mission was launched in June 2024. The methodology described here requires the GRB to have a plateau emission. Thus, SVOM’s ability to quickly slew to a GRB’s location will enable quick follow-up of afterglow emission. This will lead to more observations of the plateau, especially at X-ray wavelengths, and make the GRBs observed by SVOM compatible with our methodology, albeit with careful consideration of their parameter spaces.
[…]
5.6 Redshift inference with the web app
Our web app has various functionalities, but its main focus is inferring the redshift of GRBs without $z_{obs}$. Users have the flexibility to upload a file containing the predictors discussed in this work (see Sec. 4.2.1) or manually enter the parameters of a single GRB for redshift estimation. This interface is presented in Fig. 18. This module of the web app also allows the user to obtain MICE imputations if their data contains missing values. To do so, they select the ‘Apply MICE’ button. After providing the data, the web app initiates the back-end ML codes, the results and performance of which have been presented in this work. Upon successful execution, the user is presented with individual redshift values for each GRB in a tabular manner, which can also be downloaded as a text file. This table displays the GRB IDs along with $z_{pred}$ and its associated uncertainties. Furthermore, if a user already has their custom model, trained from modules such as Model Comparision and SuperLearner, then they can use that model for predictions. For this, they need to select the ‘Do you have modeled data?’ option. However, the users can use the model trained on our data without necessarily providing their own training set. Finally, comprehensive documentation is available on the web app in the ‘Documentation’ module to help users navigate.
Figure 18: The figure shows the homepage of the app, which is the Redshift Estimator. It also shows the panel where the user needs to upload the data file, the option for applying MICE and using custom models.
6 Discussion
The findings from Dainotti et al. (2024b) demonstrate the promise of using machine learning for redshift estimation, particularly in overcoming the limited number of GRBs with precise redshift measurements. In this work, we have expanded both the methodology and the sample sizes for the training and generalization datasets. Providing redshift predictions for the generalization set will help future population studies address the GRBFR and SFR discrepancy and other relevant issues. In the future, the $z_{pred}$ values obtained here can also be utilized to explore the characteristics of the central engine in GRB events, focusing on GRBFR, LF, and the structure of LCs. We can also use GRB LCs obtained from reconstruction, which some of us did in Dainotti et al. (2023). We demonstrated the effectiveness of this reconstruction technique in reducing the errors on plateau parameters by an average of 30%-40%.
6.1 LGRBs vs SGRBs
Contemporaneously, this research can also shed light on the distinction between various GRB classes. As mentioned previously, LGRBs and SGRBs are characterized by different properties. Thus, one has to classify GRBs as accurately as possible in order to conduct high-precision population studies. Traditionally, the division between collapsar and merger type (non-collapsar) origin has been discussed mainly in terms of the $T_{90}$ distribution for the BATSE observed events. Collapsar events are characterized by $T_{90}\geq 2s$, while non-collapsar events by $T_{90}<2s$. An updated division based on $T_{90}$ observed by Swift has been discussed in Bromberg et al. (2013), where the $T_{90}$ cut has been set to 0.8. With the advent of gravitational waves (GW) observed with the case of GRB 170817 and the consequent observation of the kilonovae, it was discussed that the merger of neutron stars was the smoking gun event to distinguish between collapsar and non-collapsar and should be accompanied by the GW and the kilonova presence. However, it has also been discussed extensively how the case of GW 170817 was peculiar and an unusually low luminosity event. Moreover, there have been two GRB cases that have properties of LGRBs in terms of $T_{90}>$2s, but there was evidence of an associated kilonovae (KN): these are the instances of GRB 211211A and GRB 230307A. The latter case has a larger value for the luminosity of the KN=$10^{42}ergs^{-1}$.
6.2 Issue of LLGRBs
This leads us to another pertinent question about the true nature and origin of LLGRBs. Research by Coward (2005); Liang et al. (2007a); Bernardini et al. (2007); Virgili et al. (2009); Patel et al. (2023) suggests that LLGRBs represent a distinct class of phenomena. Due to their low luminosities, these events are typically observed at low redshifts, such as GRB 060218 at $z=0.033$ (Mirabal et al., 2006) and GRB 980425 at $z=0.0085$ (Tinney et al., 1998). Coward (2005) simulated the local density rate of LLGRBs to be 220 Gpc-3yr-1, based on the under-luminous GRB 980425. Under the assumption that the luminosity of GRB 980425 was typical of such a population of LLGRBs, the authors claim the luminosity cut-off between LLGRBs and high-luminosity GRBs (HLGRBs) to be $2\times 10^{-53}$erg s-1. However, based on GRB 980425 and GRB 060218 Liang et al. (2007a) defined LLGRBs as GRBs with luminosity $<$ 10-49erg s-1. The authors found the LLGRB event rate to be $\sim 325^{+352}_{-177}\,\rm Gpc^{-3}yr^{-1}$ while that of HLGRBs to be $1.12^{+0.43}_{-0.2}$ Gpc-3 yr-1. They also showed that the luminosity functions of LLGRBs and HLGRBs are different and cannot be represented by a single function. Chapman et al. (2007) also investigated the local density rate of LLGRBs based on GRB 980425 and GRB 060218 and found a result of $700\pm 360$ Gpc-3 yr-1 within a distance of 155 Mpc. Virgili et al. (2009) provided further evidence that LLGRBs are a distinct class of GRBs from HLGRBs. They demonstrated that a simple power-law LF is insufficient and that a smooth broken power-law fit is necessary to obtain a good prediction for the distribution of a combined sample of LLGRBs and HLGRBs. Using MC simulation and the LF, they derive an event rate of $\sim 100-400$ Gpc-3 yr-1 for LLGRBs and $\sim 1$ Gpc-3 yr-1 for HLGRBs. However, they conclude that the sample is too small and more data is needed to draw stronger conclusions. Dainotti et al. (2024b) obtained a GRB event rate of $8.47-9$ Gpc-3 $yr^{-1}$ between $2.31.9<z<$ 2.3 using $z_{obs}$. This matches the HLGRB rate found by Dong et al. (2023); Liang et al. (2007a); Virgili et al. (2009). Lan et al. (2021) utilized 5 LLGRBs out of a collection of $1$ LGRBs obtained from Swift to demonstrate that a triple power-law LF provides a superior fit for the data, again indicating the existence of LLGRBs as a separate component. Recently, Dong et al. (2023) demonstrated that the event rate of HLGRB closely follows the SFR, while that of LLGRB shows an excess at low-$z$ ($z<1$). Combining the sample leads to the observed low-$z$ excess and mismatch between the GRBFR and SFR. These previous works strongly suggest that LLGRBs are a distinct class of GRBs with the potential to solve the GRB-SFR discrepancy. However, to draw more definitive conclusions, a larger sample of GRBs with redshift is needed.
6.3 Discrepancy with SFR
This discrepancy with the SFR could also be due to the effects of collimated emission within narrow jets not being taken into account in the current analysis.
Indeed, we need such collimated emission to achieve the observed high luminosity. Unlike most parametric forward fitting methods, which assume both no luminosity evolution and that the GRBFR should be the same as SFR, the approach of Dainotti et al. (2024b) offers the advantage of deriving the formation rate evolution non-parametrically directly from the data. Thus, we can safely bypass the assumption about the GRB rate following the SFR. This result is also shown for LGRBs by the Konus-WIND team (Tsvetkova et al., 2017) and the other 5 teams independently (Petrosian et al., 2015; Yu et al., 2015; Pescalli et al., 2016; Lloyd-Ronning et al., 2019; Yonetoku et al., 2004), using the same statistical methods adopted by Dainotti et al. (2021c). Regarding SGRBs, including those with extended emission, Dainotti et al. (2021c) found that SGRB rates also do not follow the SFR at low-$z$, even when adding a delay for SGRBs to account for the older age of the progenitors.
6.4 Impact of our research
Our research here can help address the above-mentioned issues. First, employing 276 GRBs with $z_{pred}$ will allow the community to distinguish between GRB rate and SFR and what the GRB progenitors are, specifically regarding LLGRBs. Furthermore, the web app will enable the whole community to quickly obtain $z_{pred}$ and this will enhance the above-mentioned advantages. The web app’s versatility will help it stay updated when new GRBs with $z_{obs}$ are obtained, leading to more accurate predictions.
Finally, our MC simulations and analysis of SVOM’s upcoming data paint a very promising future for GRB cosmology. Our analysis, combined with comprehensive progenitor studies (Dainotti et al., 2020; Wang et al., 2022), will provide a set of sources that can be used as reliable cosmological tools. The high-$z$ GRBs that THESEUS will explore (expected launch 2032, Amati et al. (2021)) are crucial probes for investigating the SFR at the epoch of reionization and investigate the number of Population 3 stars.
7 Summary, and Conclusion
In this study, we have built upon the methodology established by Dainotti et al. (2024b) and Dainotti et al. (2024c). We use a 20% larger training set and more robust formula estimation to predict the redshifts of 276 GRBs. In addition, we perform an extensive MC simulation to evaluate how our ML model will perform in the upcoming years. For the benefit of the community, we have a publicly available and free-to-use web app designed to provide redshift estimates for any GRB with plateau emission, along with additional functionalities.
We draw the following conclusions from this research:
  • The ML results obtained in this analysis, presented in Sec. 5.1, have similar accuracy with those presented in Dainotti et al. (2024c) but slightly lower. Specifically, we have a 10% decrease in $r$, from 0.719 to 0.644, and a 9.8% increase in RMSE, from 0.91 to 1. These changes in the metrics can be attributed to two reasons. First, the new sample of GRBs added to the training set has a higher percentage of GRBs at larger redshift, as compared to the training set in Dainotti et al. (2024c) (see Sec. 3.1). This can lead to the ML model having to adjust to the new data with a larger intrinsic scatter, thus leading to a decrease in the performance metrics. The second reason is the change in the methodology for formula estimation. In Dainotti2024GRBRedshift, the formulas estimated for GAM and GLM were determined using a single training-test split instance. Here, we have expanded the methodology and generalized the formulas. Generalizing the formula here means it can better capture the underlying relations among GRB properties for a larger population of GRBs rather than just a small, limited sample of the training set. However, this generalization leads to poorer performance on a singular training-test split as the variability of the underlying relation is higher in a smaller sample. However, the formulas generalize well to additional synthetic GRB samples with a more uniform redshift distribution, which is evident from the improved balanced sampling result across all the metrics (lower right and left panels of Fig. 13). This promising result indicates that our ML model has effectively generalized to the entire GRB sample and the MC simulations further emphasize this point.
  • In Section 5.2, we demonstrated that $\log(NH)$ exerts the greatest influence on redshift predictions. This finding is consistent with the results reported in Dainotti et al. (2024c) and Dainotti et al. (2024b). A similar result was also discovered independently by Racz (2023), where they found that the performance of their ML model was reduced when they removed $\log(NH)$ from their list of predictors. We also find that $\log(T_{a})$ is the second most important predictor of redshift, similar to Dainotti et al. (2024c), whereas Dainotti et al. (2024b) found that $\log(F_{a})$ is the second most predictive feature. Given the correlation of $\log(T_{a})$ and $\log(F_{a})$, these results demonstrate that plateau features are important when predicting the redshift of GRBs.
  • In Sec 5.3, we predict the redshift of 276 GRBs using our trained model. Both the bias-corrected and non-bias-corrected predictions are freely available to the community through this publication. However, we still have a discrepancy between the predicted redshift distributions and the redshift distribution of the training set. Dainotti et al. (2024c) questioned as to what extent the variation between the redshift distributions of GRBs with and without $z_{obs}$ are inherently different, or if the variation might be attributed to the limited size of the training set. In this analysis, we have increased the training set with 30 additional GRBs. However, we still need to investigate the origin of the different distributions. A KS-test comparing the $\log(NH)$ distributions for the training and the generalization sets still presents us with the conclusion that they are derived from different parent populations. Given the fact that $\log(NH)$ is the most important predictor, this could explain the discrepancy in the redshift distributions.
  • The MC simulations of future data in Sec. 5.4 present us with promising results. Based on those estimates, we conclude that our ML model is robust enough to last for at least three more years before major retraining is needed. The histogram distributions for $r$, RMSE, and MAD show that our ML model performs better when more sources are added to the training set. It’s important to note that this condition holds only if the additional data incorporated into the training set in the future comes from the same parent population as the existing sample.
  • The methodology described here can be accessed via the associated web app. It enables users to obtain redshift predictions based on the ML model presented here for their GRBs. It also enables users to choose from a set of predefined models that, in our experience, work well with GRB samples and train these models with their data. Then, they can obtain redshift predictions from this custom model. The web app is designed to provide results like the ones presented in this work for any custom data the user may input.
  • In Sec 5.5, we presented calculations on how many GRB’s redshift estimates we will be able to obtain in the future using SVOM data, and possibly THESEUS. The numbers mentioned there are based on current estimates from publications of SVOM and Swift’s observational statistics. These numbers will need to be updated as more GRBs by SVOM are detected. Regardless, the numbers paint a promising picture of the future applicability of such ML techniques. In fact, this work has already enabled us to increase the Swift GRB sample with redshift by $\sim$110%.
8 Acknowledgments
The authors thank Omar Elsherif for his contributions to building the basic structure of the webapp, upon which the whole functionality is based. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017205. We acknowledge Biagio de Simone and Christopher Cook for helping with the MCcomputations. We acknowledge funding from Swift GI Cycle XIX, grant number 22-SWIFT22-0032.